A scientist has a beaker contain 20 mL of a solution containing 20% acid. To dilute this pure water is added.
a. Write the equation for the concentration in the beaker after adding n mL of water.
b. Find the concentration if 10 mL of water is added
c. How much mL of water must be added to equal 4%
d. what is the behavior as n approaches infinity, and what is the physical significance of this?

Answer by stanbon(75887)

A scientist has a beaker contain 20 mL of a solution containing 20% acid. To dilute this pure water is added.
a. Write the equation for the concentration in the beaker after adding n mL of water.
Equation:
acid + acid = acid
0.20*20 + 0*n = x(20+n)
----------------------------
b. Find the concentration if 10 mL of water is added
0.20*20 + 0*10 = x(20+10)
Multiply thru by 100 to get::
20*20 + 0 = x(3000)
x = 400/(3000
x = 13.3%
----------------
c. How much mL of water must be added to equal 4%
0.20*20 + 0*n = 0.04(20+n)
20*20 + 0 = 4(20+n)
100 = 20+n
n = 80 ml
------------------
d. what is the behavior as n approaches infinity, and what is the physical significance of this?
As more and more water is added the concentration approaches 0%
=============
Cheers,
Stan H.
-----------------